# Jo Andersson Stokke 22    Ronny Bergmann 22    Martin Hanik 113344    Christoph von Tycowicz 44

### v1

@@ Abstract
Keywords:
Hypergraph Laplacian Riemannian Geometry Diffusion
@@ 2.2 Differential Geometry
[…]
@@ 3.2 Gradients and Laplacians
$\nabla^{F}f(e):=\sqrt{w(e)}\log_{x_{\mathrm{in}}}x_{\mathrm{out}}$
@@
$\nabla^{P}f(e)(u,v):=\frac{\sqrt{w(e)}}{|e_{\mathrm{in}}||e_{\mathrm{out}}|}\log_{f(u)}f(v).$
@@
    $\nabla^{F}f(e)=-\mathrm{PT}_{x_{\mathrm{out}}\rightarrow x_{\mathrm{in}}}\nabla^{F}f(\tilde{e})$
@@
$\Delta_{p}^{\mathrm{i},F}f(u)$ $:=\frac{-1}{|\mathcal{N}^{\mathrm{in}}(u)|^{\eta}}\left(\sum_{e\in\mathcal{N}^{\mathrm{in}}(u)}\frac{w(e)\,d_{\mathcal{M}}^{2}(x_{\mathrm{in}},x_{\mathrm{out}})}{|e_{\mathrm{in}}|}\right)^{\frac{p-2}{2}}$
@@
$\Delta_{p}^{\mathrm{a},F}f(u):=\frac{-1}{|\mathcal{N}^{\mathrm{in}}(u)|^{\eta}}\sum_{e\in\mathcal{N}^{\mathrm{in}}(u)}\frac{\sqrt{w(e)}^{p}\,d_{\mathcal{M}}^{p-2}(x_{\mathrm{in}},x_{\mathrm{out}})}{|e_{\mathrm{in}}|}\mathrm{PT}_{x_{\mathrm{in}}\rightarrow f(u)}\log_{x_{\mathrm{in}}}x_{\mathrm{out}}$
@@
$\Delta_{p}^{\mathrm{i},P}f(u):=$ $\frac{-1}{|\mathcal{N}^{\mathrm{in}}(u)|^{\eta}}\left(\sum_{e\in\mathcal{N}^{\mathrm{in}}(u)}\sum_{u_{1}\in e_{\mathrm{in}}}\sum_{u_{2}\in e_{\mathrm{out}}}\frac{w(e)\,d_{\mathcal{M}}^{2}(f(u_{1}),f(u_{2}))}{|e_{\mathrm{in}}|^{2}|e_{\mathrm{out}}|}\right)^{\frac{p-2}{2}}$
@@ 4 Experiment
We examine diffusion dynamics of the normalized Fréchet and pairwise $2$-Laplacian on unit-weight hypergraphs with features on the $2$-sphere $\mathcal{M}=\mathbb{S}^{2}$ with its canonical metric. (Computations on the sphere were performed with Morphomatics [1].) To this end, we randomly generate signed incidence matrices representing symmetric oriented hypergraphs and embed them on an octant of $\mathbb{S}^{2}$ by sampling angles $\theta,\phi$ from the uniform distribution on $[0,\frac{\pi}{2})\times[0,\frac{\pi}{2}]$. With $\Delta_{2}^{*}\in\{\Delta_{2}^{F},\Delta_{2}^{P}\}$, we compute equilibria of the heat diffusion equation
@@
These results exemplify that the hypergraph Laplacians exhibit a richer set of equilibria. While we expect the graph Laplacian for a positively-weighted graph embedded inside a so-called normal convex neighborhood [^2] An open neighborhood $U\subseteq\mathcal{M}$ is called normal convex if all $p,q\in U$ are connected by a geodesic that does not leave $U$.—like an octant of the sphere—to vanish only for constant configurations, we observe convergence to non-constant solutions with hypergraph.

### v2

# Jo Andersson Stokke    Ronny Bergmann    Martin Hanik    Christoph von Tycowicz

@@ 2.2 Differential Geometry
Let $(\mathcal{M},\langle\cdot,\cdot\rangle)$ be a connected, complete Riemannian manifold with Riemannian metric $\langle\cdot,\cdot\rangle$ and induced distance $d_{\mathcal{M}}:\mathcal{M}\times\mathcal{M}\to\mathbb{R}$. We denote the Riemannian exponential and logarithm at $x\in\mathcal{M}$ by $\exp_{x}$ and $\log_{x}$, respectively. Since $\mathcal{M}$ is complete, a length-minimizing geodesic segment $\gamma_{xy}$ connects the two points for every pair $x,y\in\mathcal{M}$. This segment is unique if $y$ is inside the injectivity domain of $\exp_{x}$, and, in that case, we denote the parallel transport of a vector $X\in T_{x}\mathcal{M}$ to $T_{y}\mathcal{M}$ along $\gamma_{xy}$ by $\operatorname{PT}_{x\rightarrow y}X$.
@@ 3.2 Gradients and Laplacians
$\nabla^{F}f(e)=\sqrt{w(e)}\log_{x_{\mathrm{in}}}x_{\mathrm{out}}$
@@
$\nabla^{P}f(e)(u,v)=\frac{\sqrt{w(e)}}{|e_{\mathrm{in}}||e_{\mathrm{out}}|}\log_{f(u)}f(v).$
@@
[…]
@@
$\Delta_{p}^{\mathrm{i},F}f(u)$ $=\frac{-1}{|\mathcal{N}^{\mathrm{in}}(u)|^{\eta}}\left(\sum_{e\in\mathcal{N}^{\mathrm{in}}(u)}\frac{w(e)\,d_{\mathcal{M}}^{2}(x_{\mathrm{in}},x_{\mathrm{out}})}{|e_{\mathrm{in}}|}\right)^{\frac{p-2}{2}}$
@@
$\Delta_{p}^{\mathrm{a},F}f(u)=\frac{-1}{|\mathcal{N}^{\mathrm{in}}(u)|^{\eta}}\sum_{e\in\mathcal{N}^{\mathrm{in}}(u)}\frac{\sqrt{w(e)}^{p}\,d_{\mathcal{M}}^{p-2}(x_{\mathrm{in}},x_{\mathrm{out}})}{|e_{\mathrm{in}}|}\mathrm{PT}_{x_{\mathrm{in}}\rightarrow f(u)}\log_{x_{\mathrm{in}}}x_{\mathrm{out}}$
@@
$\Delta_{p}^{\mathrm{i},P}f(u)$= $\frac{-1}{|\mathcal{N}^{\mathrm{in}}(u)|^{\eta}}\left(\sum_{e\in\mathcal{N}^{\mathrm{in}}(u)}\sum_{u_{1}\in e_{\mathrm{in}}}\sum_{u_{2}\in e_{\mathrm{out}}}\frac{w(e)\,d_{\mathcal{M}}^{2}(f(u_{1}),f(u_{2}))}{|e_{\mathrm{in}}|^{2}|e_{\mathrm{out}}|}\right)^{\frac{p-2}{2}}$
@@ 4 Experiment
We examine diffusion dynamics of the normalized Fréchet and pairwise $2$-Laplacian on unit-weight hypergraphs with features on the $2$-sphere $\mathcal{M}=\mathbb{S}^{2}$ with its canonical metric. (Computations on the sphere were performed with Morphomatics [1].) To this end, we randomly generate signed incidence matrices representing symmetric oriented hypergraphs and embed them on an octant of $\mathbb{S}^{2}$ by sampling angles $\theta,\phi$ from the uniform distribution on $0,\frac{\pi}{2})\times[0,\frac{\pi}{2}]$. With $\Delta_{2}^{*}\in\{\Delta_{2}^{F},\Delta_{2}^{P}\}$, we compute equilibria of the heat diffusion equation
@@
These results exemplify that the hypergraph Laplacians exhibit a richer set of equilibria. While we expect the graph Laplacian for a positively-weighted graph embedded inside a so-called normal convex neighborhood [^2] An open neighborhood $U\subseteq\mathcal{M}$ is called normal convex if all $p,q\in U$ are connected by a geodesic that does not leave UU.—like an octant of the sphere—to vanish only for constant configurations, we observe convergence to non-constant solutions with hypergraph.
@@ 5 Conclusion
5.0.1 Acknowledgments.
Funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) under Germany´s Excellence Strategy—The Berlin Mathematics Research Center MATH+ (EXC-2046/1, EXC-2046/2, project ID: 390685689).